Alex has twice as much money as Jennifer. Jennifer has $6 less than Shannon. Togethere they have $54. How much money does each have?
Answer by stanbon(75887)

Alex has twice as much money as Jennifer. Jennifer has $6 less than Shannon. Togethere they have $54. How much money does each have?
----
Equations:
a = 2j
j = s - 6
a + j + s = 54
----
Substitute for a and "s" and solve for "j:
2j + j + j+6 = 54
4j = 48
jennifer has $12
-----
a = 2j
Alex has 2*12 = $24
------
s = j+6 = 12+6
Shannon has $18
===
Cheers,
Stan H.